OpenStudy (anonymous):

What is the common denominator of 2/9 and 5/8?

OpenStudy (anonymous):

no less than 72

OpenStudy (anonymous):

Ok, I'm learning fractions and I have to add them. Do I just add them without reducing since they can't be reduced?

OpenStudy (anonymous):

no : 2*8/72 + 5*9/72 = 16/72 + 45/72 = 61/72

OpenStudy (anonymous):

Thank you